# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - Blockchain and Applications

### v2

@@ 4.3 Cryptographic Hash Functions
Cryptographic hash functions are mathematical algorithms that transform an arbitrary amount of input data into a fixed-size output, known as a hash or digest. These functions are essential in ensuring data integrity and are widely used in various applications like digital signatures, message authentication codes, and blockchain technologies[98]. In this section, we will discuss two popular cryptographic hash functions[99]: MD5 and the SHA family of hash functions.
@@ 4.3.1 MD5 Algorithm
MD5 (Message Digest Algorithm 5)[100] is one of the earliest widely-used cryptographic hash functions. It was designed by Ronald Rivest[101] in 1991 to produce a 128-bit hash value from an arbitrary-length input.
@@
The MD5 algorithm[102] processes data in 512-bit chunks, breaking it down into a series of operations over multiple rounds. The final output of MD5 is a 128-bit hash, often represented as a 32-character hexadecimal string. Here’s a high-level breakdown of how MD5 works:
@@
MD5 was widely used in the past, but its security has been severely compromised over time due to its vulnerability to collision attacks[103]. A collision occurs when two different inputs produce the same hash. In cryptographic applications, this can be catastrophic, leading to attacks such as creating fraudulent digital certificates.
@@
  • Collision vulnerability: In 2004, researchers demonstrated a practical method for finding MD5 collisions[104], making it unsuitable for applications requiring high levels of security, such as digital signatures or SSL certificates.
@@ 4.3.2 SHA Family of Hash Functions
The SHA (Secure Hash Algorithm) family includes several cryptographic hash functions designed by the National Security Agency (NSA)[105, 106]. The most common variants are SHA-1, SHA-2, and SHA-3. SHA-1 has been deprecated, while the SHA-2 family (SHA-224, SHA-256, SHA-512) is widely used today due to its improved security properties[107].
@@
SHA-1, developed in 1993, produces a 160-bit hash from input data. Its design is similar to MD5, but it processes data in 512-bit chunks, iterating over 80 rounds. SHA-1 was once widely used in digital signatures, SSL certificates, and other security protocols. However, it was officially deprecated by NIST in 2011[108] due to its vulnerability to collision attacks.
@@
  • Collision attacks: In 2017, researchers at Google and CWI Amsterdam successfully demonstrated a collision attack on SHA-1[109], making it insecure for most cryptographic applications.
@@ 4.3.3 Detailed Analysis of SHA-256 Algorithm
SHA-256 (Secure Hash Algorithm 256) is a member of the SHA-2 family of cryptographic hash functions[110]. It produces a fixed-size 256-bit (32-byte) hash value, often referred to as a message digest. SHA-256 is widely used in cryptography and blockchain technology, especially for ensuring the integrity of data.
@@ 4. Compression Function
A Merkle Tree is a binary tree structure used to efficiently verify the integrity of large datasets[111]. In blockchain technology, Merkle Trees are used to ensure the integrity of transactions in blocks. SHA-256 plays a crucial role in constructing Merkle Trees by generating hash values for the data at each level of the tree.
@@ 1. Bitcoin Mining
In Bitcoin, miners compete to find a value (called a nonce) such that the SHA-256 hash of the block header starts with a certain number of leading zeros[112]. This process, known as Proof of Work (PoW), requires extensive computational effort but ensures the security of the network.
@@ 2. Digital Signatures
SHA-256 is often used to create a digest of a message, which is then signed with a private key to create a digital signature[113]. This process ensures both the integrity and authenticity of the message.
@@ 3. Secure Communications
SHA-256 is used in various protocols such as TLS (Transport Layer Security) to ensure that data transmitted over the internet is secure and has not been tampered with[114]. It is also used in file integrity verification and software distribution to ensure that files have not been altered.
@@ 4.3.4 RIPEMD160 and Bitcoin Address Generation
RIPEMD160 (RACE Integrity Primitives Evaluation Message Digest)[115] is a 160-bit cryptographic hash function that plays an important role in Bitcoin address generation. Bitcoin addresses are shorter than the 256-bit public keys used in Bitcoin transactions, and this is achieved by combining the SHA-256 and RIPEMD160 hash functions.
@@
This process ensures that Bitcoin addresses are both compact and secure. By combining the outputs of SHA-256 and RIPEMD160, Bitcoin takes advantage of the security features of both algorithms while keeping the addresses manageable in length[116].
@@ 4.4.1 Ensuring Blockchain Data Integrity
Hash functions are fundamental to the security and integrity of blockchain systems. A hash function is a mathematical algorithm that transforms an input (or ”message”) into a fixed-size string of bytes. The output is commonly referred to as the hash value, digest, or simply the hash. In the context of blockchain, a specific hash function like SHA-256 (Secure Hash Algorithm)[117] is often used.
@@ 4.4.2 Application of Hash Functions in Proof-of-Work
In a blockchain network, the Proof-of-Work (PoW) consensus mechanism is crucial for verifying transactions and adding new blocks to the chain[118]. Hash functions, particularly SHA-256 in Bitcoin’s case, play a central role in this process. PoW involves solving a complex mathematical puzzle, where miners compete to find a valid hash that is below a specific target value.
@@ 4.4.3 Combining Digital Signatures with Hash Functions
In blockchain systems, digital signatures are often used alongside hash functions to guarantee the authenticity and integrity of transactions[119]. When a user initiates a transaction, the data of the transaction is first hashed. This hash is then signed using the user’s private key, creating a unique digital signature.
@@ 5.1.1 Early Electronic Money and Its Limitations
In the early days of digital communication, there were various attempts to create electronic forms of money. These early systems aimed to digitize money and make online transactions more efficient. Two prominent examples of early digital currency include eCash and DigiCash, which were proposed by cryptographer David Chaum[120] in the 1980s and 1990s.
eCash was an anonymous electronic cash system that allowed users to make online payments without revealing their identity, making it a precursor to modern privacy-focused digital currencies[121]. DigiCash, the company that developed eCash, aimed to ensure privacy in digital transactions by using cryptographic protocols that would blind the issuer (i.e., the bank) to the details of the transaction[122]. However, eCash and DigiCash faced several challenges that ultimately prevented their widespread adoption.
@@
  • Double-Spending Problem: One of the biggest challenges with digital currency is the risk of double spending, where the same digital token is spent more than once. In traditional banking systems, this issue is handled by central authorities who maintain a ledger of all transactions. However, in the absence of such authorities, it became difficult to prevent users from replicating digital tokens and spending them multiple times[123].
@@ 5.1.2 The Emergence of Bitcoin: Solving the Double-Spending Problem
Bitcoin was introduced in 2008 by an individual or group of individuals using the pseudonym Satoshi Nakamoto[124]. What made Bitcoin revolutionary was its ability to solve the double-spending problem without relying on a central authority. Prior to Bitcoin, the double-spending problem was a fundamental obstacle that digital currencies faced, as digital tokens can be easily copied, unlike physical currencies.
Bitcoin addressed this issue using a decentralized peer-to-peer network and an innovative technology called the blockchain[125]. The blockchain is a distributed ledger that records all transactions in a transparent and immutable manner. This means that once a transaction is added to the blockchain, it cannot be altered or deleted, ensuring that the same Bitcoin cannot be spent more than once.
@@
Bitcoin’s unique design, combining decentralization, blockchain technology, and cryptographic security, offered a robust and scalable solution to the double-spending problem, which had plagued earlier digital currency systems[126].
@@ 5.2.1 The Core Idea of Bitcoin: Decentralization
The core idea behind Bitcoin, as outlined in Satoshi Nakamoto’s whitepaper titled ”Bitcoin: A Peer-to-Peer Electronic Cash System,” [126] is the concept of decentralization. Unlike traditional financial systems that rely on intermediaries such as banks, Bitcoin removes the need for a trusted third party to validate transactions.
@@ 5.2.2 Satoshi Nakamoto’s Design Goals
Satoshi Nakamoto’s design for Bitcoin was shaped by several key goals, which are reflected in the architecture of the Bitcoin network[126]:
@@ 5.3.1 Proof-of-Work Mechanism
The Proof-of-Work (PoW) mechanism is fundamental to ensuring the security and integrity of the Bitcoin network[127]. PoW works by requiring participants, known as miners, to solve complex cryptographic puzzles in order to validate transactions and add new blocks to the blockchain. This process is computationally intensive, meaning that it requires significant computational power and energy.
@@ 5.3.2 Chain-Based Block Structure
Bitcoin’s blockchain structure is an innovative way to ensure data integrity and immutability. Each block in the Bitcoin blockchain contains a list of transactions, and every block is cryptographically linked to the previous one by including the hash of the prior block within its header[128]. This creates a continuous chain of blocks (hence the term blockchain), forming an immutable ledger of all transactions since the network’s inception.
@@ 5.3.3 Peer-to-Peer Network and Distributed Consensus
Bitcoin operates on a decentralized, peer-to-peer (P2P) network[129]. This means that no single entity, such as a central bank or government, controls the system. Instead, every participant (node) in the network has equal access and authority. Nodes work together to validate and propagate transactions across the network, creating a trustless environment.
Decentralization is achieved through the process of distributed consensus, where nodes independently validate new blocks and transactions by adhering to the rules set by the Bitcoin protocol[130]. Each node holds a copy of the entire blockchain, which they use to verify incoming transactions.
@@ 5.4.1 Creating a Bitcoin Wallet Using Tools
A Bitcoin wallet is a software application or hardware device that stores the private keys you need to access and spend your Bitcoin[131]. There are several types of wallets, each offering different levels of security, usability, and features. Below, we will guide you through selecting a wallet and the steps required to create one.
@@
  • Desktop Wallets: These are software applications that you download and install on your computer. Desktop wallets offer a good balance between security and convenience but are vulnerable to malware if your computer is compromised. Examples include Electrum[132] and Bitcoin Core[133].
  • Mobile Wallets: These wallets are designed for smartphones and provide a convenient way to manage your Bitcoin on the go. However, they are less secure than hardware or desktop wallets. Popular mobile wallets include Mycelium[134] and Breadwallet[135].
  • Hardware Wallets: These are physical devices designed to store your private keys offline, providing maximum security against hacking. Examples include Ledger Nano S[136] and Trezor[137].
@@
When creating a Bitcoin wallet, the wallet software generates a mnemonic phrase (sometimes called a seed phrase), which is a set of 12, 18, or 24 randomly generated words[138]. This mnemonic phrase is a human-readable representation of the private key, and it can be used to recover the wallet if the device where the wallet was created is lost or destroyed.
@@ 5.4.2 How Bitcoin Wallets Work
Bitcoin wallets rely on Elliptic Curve Cryptography (ECC) for generating public and private key pairs[139]. ECC is a powerful cryptographic method that provides a high level of security with relatively small key sizes. In the case of Bitcoin, ECC operates over a specific curve called secp256k1. The curve allows for secure key generation and ensures that it is computationally infeasible to derive the private key from the public key.
@@
The elliptic curve is a mathematical structure that enables the generation of keys[140]. Each point on the curve corresponds to a valid key pair. The private key is simply a randomly chosen large number, while the public key is a point on the elliptic curve derived from the private key using a multiplication operation. This one-way function ensures that while it is easy to derive the public key from the private key, it is nearly impossible to reverse the process.
@@
The public key is derived from the private key using an elliptic curve multiplication operation. In Bitcoin, this multiplication is done over the secp256k1 curve[141].
@@ 6.1.1 What is Blockchain?
Blockchain is a decentralized, distributed ledger technology that allows data (typically transactions) to be recorded and maintained across many computers, known as nodes[142]. Unlike traditional centralized systems where a central authority (such as a bank) manages records, blockchain operates without any single entity in control. Instead, it is managed by a network of participants, each of whom holds a copy of the entire ledger.
@@
For example, if a person wants to send cryptocurrency to another, a transaction will be created, broadcast to the network, and validated through the consensus mechanism[143]. Once validated, the transaction is grouped with others into a block, and this block is added to the existing chain of blocks.

### v3

@@ 4.3 Cryptographic Hash Functions
Cryptographic hash functions are mathematical algorithms that transform an arbitrary amount of input data into a fixed-size output, known as a hash or digest. These functions are essential in ensuring data integrity and are widely used in various applications like digital signatures, message authentication codes, and blockchain technologies[Dang2009RecommendationFA]. In this section, we will discuss two popular cryptographic hash functions[Pethe2016AnOO]: MD5 and the SHA family of hash functions.
@@ 4.3.1 MD5 Algorithm
MD5 (Message Digest Algorithm 5)[Wang2005HowTB] is one of the earliest widely-used cryptographic hash functions. It was designed by Ronald Rivest[Wang2004CollisionsFH] in 1991 to produce a 128-bit hash value from an arbitrary-length input.
@@
The MD5 algorithm[Wartik1992HashingA] processes data in 512-bit chunks, breaking it down into a series of operations over multiple rounds. The final output of MD5 is a 128-bit hash, often represented as a 32-character hexadecimal string. Here’s a high-level breakdown of how MD5 works:
@@
MD5 was widely used in the past, but its security has been severely compromised over time due to its vulnerability to collision attacks[Stevens2012AttacksOH]. A collision occurs when two different inputs produce the same hash. In cryptographic applications, this can be catastrophic, leading to attacks such as creating fraudulent digital certificates.
@@
  • Collision vulnerability: In 2004, researchers demonstrated a practical method for finding MD5 collisions[Mikle2004PracticalAO], making it unsuitable for applications requiring high levels of security, such as digital signatures or SSL certificates.
@@ 4.3.2 SHA Family of Hash Functions
The SHA (Secure Hash Algorithm) family includes several cryptographic hash functions designed by the National Security Agency (NSA)[Shah2014SECUREDHA, Khan2022EVOLUTIONAA]. The most common variants are SHA-1, SHA-2, and SHA-3. SHA-1 has been deprecated, while the SHA-2 family (SHA-224, SHA-256, SHA-512) is widely used today due to its improved security properties[Eastlake2006USSH].
@@
SHA-1, developed in 1993, produces a 160-bit hash from input data. Its design is similar to MD5, but it processes data in 512-bit chunks, iterating over 80 rounds. SHA-1 was once widely used in digital signatures, SSL certificates, and other security protocols. However, it was officially deprecated by NIST in 2011[MAlNawashi2024AnalysisAE] due to its vulnerability to collision attacks.
@@
  • Collision attacks: In 2017, researchers at Google and CWI Amsterdam successfully demonstrated a collision attack on SHA-1[Stevens2017TheFC], making it insecure for most cryptographic applications.
@@ 4.3.3 Detailed Analysis of SHA-256 Algorithm
SHA-256 (Secure Hash Algorithm 256) is a member of the SHA-2 family of cryptographic hash functions[2019ASO]. It produces a fixed-size 256-bit (32-byte) hash value, often referred to as a message digest. SHA-256 is widely used in cryptography and blockchain technology, especially for ensuring the integrity of data.
@@ 4. Compression Function
A Merkle Tree is a binary tree structure used to efficiently verify the integrity of large datasets[Jing2021ReviewAI]. In blockchain technology, Merkle Trees are used to ensure the integrity of transactions in blocks. SHA-256 plays a crucial role in constructing Merkle Trees by generating hash values for the data at each level of the tree.
@@ 1. Bitcoin Mining
In Bitcoin, miners compete to find a value (called a nonce) such that the SHA-256 hash of the block header starts with a certain number of leading zeros[Chow2017TheBM]. This process, known as Proof of Work (PoW), requires extensive computational effort but ensures the security of the network.
@@ 2. Digital Signatures
SHA-256 is often used to create a digest of a message, which is then signed with a private key to create a digital signature[Radack2009TheCH]. This process ensures both the integrity and authenticity of the message.
@@ 3. Secure Communications
SHA-256 is used in various protocols such as TLS (Transport Layer Security) to ensure that data transmitted over the internet is secure and has not been tampered with[Turner2014TransportLS]. It is also used in file integrity verification and software distribution to ensure that files have not been altered.
@@ 4.3.4 RIPEMD160 and Bitcoin Address Generation
RIPEMD160 (RACE Integrity Primitives Evaluation Message Digest)[Preneel1TC] is a 160-bit cryptographic hash function that plays an important role in Bitcoin address generation. Bitcoin addresses are shorter than the 256-bit public keys used in Bitcoin transactions, and this is achieved by combining the SHA-256 and RIPEMD160 hash functions.
@@
This process ensures that Bitcoin addresses are both compact and secure. By combining the outputs of SHA-256 and RIPEMD160, Bitcoin takes advantage of the security features of both algorithms while keeping the addresses manageable in length[Li2023NewRI].
@@ 4.4.1 Ensuring Blockchain Data Integrity
Hash functions are fundamental to the security and integrity of blockchain systems. A hash function is a mathematical algorithm that transforms an input (or ”message”) into a fixed-size string of bytes. The output is commonly referred to as the hash value, digest, or simply the hash. In the context of blockchain, a specific hash function like SHA-256 (Secure Hash Algorithm)[Sharma2023DifferentCH] is often used.
@@ 4.4.2 Application of Hash Functions in Proof-of-Work
In a blockchain network, the Proof-of-Work (PoW) consensus mechanism is crucial for verifying transactions and adding new blocks to the chain[Patil2023UnderstandingCM]. Hash functions, particularly SHA-256 in Bitcoin’s case, play a central role in this process. PoW involves solving a complex mathematical puzzle, where miners compete to find a valid hash that is below a specific target value.
@@ 4.4.3 Combining Digital Signatures with Hash Functions
In blockchain systems, digital signatures are often used alongside hash functions to guarantee the authenticity and integrity of transactions[Zhang2019SecurityAP]. When a user initiates a transaction, the data of the transaction is first hashed. This hash is then signed using the user’s private key, creating a unique digital signature.
@@ 5.1.1 Early Electronic Money and Its Limitations
In the early days of digital communication, there were various attempts to create electronic forms of money. These early systems aimed to digitize money and make online transactions more efficient. Two prominent examples of early digital currency include eCash and DigiCash, which were proposed by cryptographer David Chaum[Moore2013ThePA] in the 1980s and 1990s.
eCash was an anonymous electronic cash system that allowed users to make online payments without revealing their identity, making it a precursor to modern privacy-focused digital currencies[Gouget2008RecentAI]. DigiCash, the company that developed eCash, aimed to ensure privacy in digital transactions by using cryptographic protocols that would blind the issuer (i.e., the bank) to the details of the transaction[chaum2015beginnings]. However, eCash and DigiCash faced several challenges that ultimately prevented their widespread adoption.
@@
  • Double-Spending Problem: One of the biggest challenges with digital currency is the risk of double spending, where the same digital token is spent more than once. In traditional banking systems, this issue is handled by central authorities who maintain a ledger of all transactions. However, in the absence of such authorities, it became difficult to prevent users from replicating digital tokens and spending them multiple times[El2020BitcoinCI].
@@ 5.1.2 The Emergence of Bitcoin: Solving the Double-Spending Problem
Bitcoin was introduced in 2008 by an individual or group of individuals using the pseudonym Satoshi Nakamoto[Meiklejohn2018TheLO]. What made Bitcoin revolutionary was its ability to solve the double-spending problem without relying on a central authority. Prior to Bitcoin, the double-spending problem was a fundamental obstacle that digital currencies faced, as digital tokens can be easily copied, unlike physical currencies.
Bitcoin addressed this issue using a decentralized peer-to-peer network and an innovative technology called the blockchain[Lee2016NewKO]. The blockchain is a distributed ledger that records all transactions in a transparent and immutable manner. This means that once a transaction is added to the blockchain, it cannot be altered or deleted, ensuring that the same Bitcoin cannot be spent more than once.
@@
Bitcoin’s unique design, combining decentralization, blockchain technology, and cryptographic security, offered a robust and scalable solution to the double-spending problem, which had plagued earlier digital currency systems[Hunt2008BitcoinAP].
@@ 5.2.1 The Core Idea of Bitcoin: Decentralization
The core idea behind Bitcoin, as outlined in Satoshi Nakamoto’s whitepaper titled ”Bitcoin: A Peer-to-Peer Electronic Cash System,” [Hunt2008BitcoinAP] is the concept of decentralization. Unlike traditional financial systems that rely on intermediaries such as banks, Bitcoin removes the need for a trusted third party to validate transactions.
@@ 5.2.2 Satoshi Nakamoto’s Design Goals
Satoshi Nakamoto’s design for Bitcoin was shaped by several key goals, which are reflected in the architecture of the Bitcoin network[Hunt2008BitcoinAP]:
@@ 5.3.1 Proof-of-Work Mechanism
The Proof-of-Work (PoW) mechanism is fundamental to ensuring the security and integrity of the Bitcoin network[Garay2017ProofsOW]. PoW works by requiring participants, known as miners, to solve complex cryptographic puzzles in order to validate transactions and add new blocks to the blockchain. This process is computationally intensive, meaning that it requires significant computational power and energy.
@@ 5.3.2 Chain-Based Block Structure
Bitcoin’s blockchain structure is an innovative way to ensure data integrity and immutability. Each block in the Bitcoin blockchain contains a list of transactions, and every block is cryptographically linked to the previous one by including the hash of the prior block within its header[Senthilkumar2021DataCI]. This creates a continuous chain of blocks (hence the term blockchain), forming an immutable ledger of all transactions since the network’s inception.
@@ 5.3.3 Peer-to-Peer Network and Distributed Consensus
Bitcoin operates on a decentralized, peer-to-peer (P2P) network[Donet2014TheBP]. This means that no single entity, such as a central bank or government, controls the system. Instead, every participant (node) in the network has equal access and authority. Nodes work together to validate and propagate transactions across the network, creating a trustless environment.
Decentralization is achieved through the process of distributed consensus, where nodes independently validate new blocks and transactions by adhering to the rules set by the Bitcoin protocol[Jayabalan2021ASO]. Each node holds a copy of the entire blockchain, which they use to verify incoming transactions.
@@ 5.4.1 Creating a Bitcoin Wallet Using Tools
A Bitcoin wallet is a software application or hardware device that stores the private keys you need to access and spend your Bitcoin[Vishawjyoti2021EWallet]. There are several types of wallets, each offering different levels of security, usability, and features. Below, we will guide you through selecting a wallet and the steps required to create one.
@@
  • Desktop Wallets: These are software applications that you download and install on your computer. Desktop wallets offer a good balance between security and convenience but are vulnerable to malware if your computer is compromised. Examples include Electrum[electrumElectrumBitcoin] and Bitcoin Core[bitcoinBitcoinCore].
  • Mobile Wallets: These wallets are designed for smartphones and provide a convenient way to manage your Bitcoin on the go. However, they are less secure than hardware or desktop wallets. Popular mobile wallets include Mycelium[myceliumMyceliumBitcoin] and Breadwallet[breadwallet].
  • Hardware Wallets: These are physical devices designed to store your private keys offline, providing maximum security against hacking. Examples include Ledger Nano S[ledgerHardwareWallet] and Trezor[trezorTrezorHardware].
@@
When creating a Bitcoin wallet, the wallet software generates a mnemonic phrase (sometimes called a seed phrase), which is a set of 12, 18, or 24 randomly generated words[Sans2022ADM]. This mnemonic phrase is a human-readable representation of the private key, and it can be used to recover the wallet if the device where the wallet was created is lost or destroyed.
@@ 5.4.2 How Bitcoin Wallets Work
Bitcoin wallets rely on Elliptic Curve Cryptography (ECC) for generating public and private key pairs[Bolfing2020Bitcoin]. ECC is a powerful cryptographic method that provides a high level of security with relatively small key sizes. In the case of Bitcoin, ECC operates over a specific curve called secp256k1. The curve allows for secure key generation and ensures that it is computationally infeasible to derive the private key from the public key.
@@
The elliptic curve is a mathematical structure that enables the generation of keys[Yan2022TheOO]. Each point on the curve corresponds to a valid key pair. The private key is simply a randomly chosen large number, while the public key is a point on the elliptic curve derived from the private key using a multiplication operation. This one-way function ensures that while it is easy to derive the public key from the private key, it is nearly impossible to reverse the process.
@@
The public key is derived from the private key using an elliptic curve multiplication operation. In Bitcoin, this multiplication is done over the secp256k1 curve[Balasubramanian2024SecurityOT].
@@ 6.1.1 What is Blockchain?
Blockchain is a decentralized, distributed ledger technology that allows data (typically transactions) to be recorded and maintained across many computers, known as nodes[Gutlapalli2016CommercialAO]. Unlike traditional centralized systems where a central authority (such as a bank) manages records, blockchain operates without any single entity in control. Instead, it is managed by a network of participants, each of whom holds a copy of the entire ledger.
@@
For example, if a person wants to send cryptocurrency to another, a transaction will be created, broadcast to the network, and validated through the consensus mechanism[Peng2023ARO]. Once validated, the transaction is grouped with others into a block, and this block is added to the existing chain of blocks.